OpenStudy (anonymous):

$120,000 was paid to of Washburn's salespeoples in override commissions(a)what was the average amount paid to each salesperson?

OpenStudy (anonymous):

How many people?

OpenStudy (anonymous):

16

OpenStudy (anonymous):

120 000 / 16 = 7500$

OpenStudy (anonymous):

it would be $120,000 divided by the number of salespeople if it's 16 people: 120,000 divided by 16 = 7500

OpenStudy (anonymous):

120 000/16 = 7500 Assuming that money is being split equally, you would divide by 16 because that's the definition of division. If there were specific numbers given to specific people, then you would do something quite different.

OpenStudy (anonymous):

thank u so much

OpenStudy (anonymous):

Actually bluepig148 is wrong, since the question specifies that you're trying to find the average value handed out. Even if someone has received 120000$ and everyone else had received 0$, the average would still be 120000/16 = 7500.

OpenStudy (anonymous):

Right, my bad.